How do I find delta H?
II. Heat of dissolution of ammonium nitrate: Data and calculations Mass of water in cup Final temp. of water +NH,NO, Initial temp. of water AT Water 30 19 21° 2. AH Dissolution NH,NO3-
Answer #1

Mass of water (m) = 30 g

Heat capacity of water (c) = 4.184 J/g.oC

Change in temperature = Final - Initial = 19 - 21 = -2oC

Heat change = m*c* change in temperature = 30 g *4.184 J/g.oC( -2oC) = -251.04 J

Negative sign shows the heat loss by water

Heat loss by water = Heat gain by water

So, \DeltaH dissolution NH4NO3 = 251.04 J

Mass of NH4NO3 dissolved = 1.7 g

Molar Mass of NH4NO3 =80.043 g/mol

Moles of NH4NO3= Mass/ Molar mass = 1.7 g/ (80.043 g/mol) = 0.0212 mol

\DeltaH = 251.04 J /0.0212 mol = 11819.997 J/mol = 11.82 kJ/mol